# Differentially Private Sampling via Reveal-or-Obscure

### v2

@@ Abstract
We introduce a differentially private (DP) algorithm called Reveal-or-Obscure (ROO) to generate a single representative sample from a dataset of $n$ i.i.d. observations from an unknown distribution. Unlike methods that add explicit noise to the estimated empirical distribution, ROO achieves $\epsilon$-differential privacy by choosing whether to “reveal" or “obscure" the empirical distribution with a fixed probability $q$. While our proposed mechanism is structurally identical to an algorithm proposed by Cheu and Nayak, we prove a strictly better bound on the sampling complexity than that established in their theorem. Building on this framework, we propose a novel generalized sampler called Data-Specific ROO (DS-ROO), where the obscuring probability $q$ is a function of the empirical distribution. We show that when the dataset contains enough samples from every element of the alphabet, DS-ROO can achieve $\epsilon$-DP while obscuring much less. In addition, we provide tight upper bounds on the utility of DS-ROO in terms of total variation distance. Our results show that under the same privacy budget, DS-ROO can achieve better utility than state-of-the-art private samplers and vanilla ROO, with total variation distance decaying exponentially in dataset size $n$.
@@ I Introduction
The task of synthetic data generation is closely related to the broader problem of learning probability distributions, which has been studied under both central and local DP; for example, see [4, 5, 6, 7]. In the non-private setting, a learning algorithm approximates a distribution from which one can sample new data points that are representative of the original data. When privacy constraints are introduced, learning a distribution becomes significantly more challenging. In many practical cases, it may be sufficient to produce a small number of representative samples instead.
The task of privately releasing one sample—known as DP sampling—is easier than full-fledged learning, since it requires less information from the underlying distribution. Motivated by this, we introduce a novel DP sampling algorithm for discrete distributions on a finite alphabet. The key idea of our approach is to “obscure” the empirical distribution of the input dataset with a certain probability, or “reveal” it otherwise. Hence, we call our proposed algorithm reveal-or-obscure (ROO). ROO can be viewed as a form of randomized response, where it outputs a sample drawn from a mixture of two distributions. We further generalize this idea through a data-specific algorithm called Data-Specific ROO (DS-ROO), in which the probability of sampling from the uniform distribution is a function of the smallest empirical probability of the dataset. This adaptive design enables DS-ROO to obscure the empirical distribution less often when the dataset contains sufficiently many samples of each element, leading to significantly improved accuracy under the same privacy budget. Table I summarizes the total variation distance achieved by ROO, DS-ROO, and several existing private samplers, illustrating the utility gains obtained by our proposed mechanisms.
Main Contributions. We focus on the setting where we are given a dataset of $n$ samples drawn from an unknown distribution with support size $k$ and a privacy parameter $\epsilon$. In keeping with prior work, we define accuracy (utility) via the total variation distance. Our main contributions are as follows:
@@
  • We propose ROO—a sampling algorithm that achieves differential privacy without explicitly perturbing the empirical distribution of the input dataset. We incorporate uncertainty in our algorithm by sampling from the uniform distribution with some fixed probability $q$, which we henceforth refer to as obscuring probability.
@@
  • We prove that our proposed algorithm ROO achieves $\epsilon$-DP and $\alpha$-accuracy. Additionally, we show that ROO reduces the sampling complexity while achieving better privacy-utility trade-off than the state-of-the art [8, 9].
@@
  • We also propose Data-Specific ROO (DS-ROO) as a technique to generalize ROO by making $q$, i.e., the probability of sampling from the uniform distribution, a function of the empirical distribution of the dataset. We prove that DS-ROO achieves $\epsilon$-DP and its utility is always at least as good as ROO’s. In other words, the data-specific obscuring probability is no larger than that of the vanilla ROO algorithm.
  • We provide theoretical upper bounds on the utility of DS-ROO, quantified by its total variation distance, under the assumption that each element of the alphabet appears in the true distribution with probability at least $\gamma$. In particular, we show that the total variation distance goes exponentially to zero as dataset size $n\to\infty$. Our bound on the utility of DS-ROO, asymptotic in $n$ as $n\to\infty$, is shown in Table I.
@@
  • We demonstrate numerically that, for the same privacy guarantee, DS-ROO achieves significantly better utility than vanilla ROO as well as the state-of-the-art methods [8, 9].
TABLE I: Comparison of Accuracy of Private Sampling Mechanisms in Terms of Total Variation Distance with Parameters: $n$ (sample size), $k$ (alphabet size), $\epsilon$ (privacy budget), $\gamma$ (smallest probability of occurrence of an element in the true distribution). The $d_{TV}$ bounds hold for all values of $n$ except for the result for DS-ROO, which holds asymptotically as $n\to\infty$.
Raskhodnikova et al. [8] Cheu & Nayak [9] ROO (Ours) [1] DS-ROO (Ours)
$d_{\mathrm{TV}}$ $\dfrac{2k}{n\epsilon}$ $\dfrac{k-1}{n\epsilon+k-1}$ $\dfrac{1-\dfrac{1}{k}}{1+\dfrac{n}{k}(e^{\epsilon}-1)}$ $\dfrac{(k-1)e^{-2n\gamma^{2}+4\gamma m_{0}+o(1)}}{1+\dfrac{n}{k}(e^{\epsilon}-1)},\text{ where}$
$m_{0}=\left\lceil\max\left\{1+\frac{1}{\epsilon}\log\left(\frac{e}{\epsilon}-e+1\right),\frac{2}{e^{\epsilon}-1}\right\}\right\rceil$
@@
Related Work. The problem of differentially private sampling from unknown distributions is first investigated in [8]. Raskhodnikova et al. [8] provide the first known bounds with $(\epsilon,\delta)$-DP guarantees on the complexity of sampling from arbitrary distributions over a discrete alphabet. DP algorithms for sampling from higher dimensional distributions such as multivariate Gaussians are presented in [10]. Husain et al. [11] consider DP sampling in the local setting, where the central aggregator cannot be trusted and each user must produce a single data record privately. Private sampling has also been studied in the local model; recent works propose minimax-optimal mechanisms for all $f$-divergences [12], and leverage public data to improve utility under local DP [13]. A key focus of these efforts has been to reduce the sample complexity of DP-assured private sampling. A recent and concurrent work of Cheu and Nayak [9] independently proposes an algorithm structurally identical to ROO. However, our analysis establishes a strictly better sampling complexity bound in the same setting.
More generally, the problem of releasing a dataset in a differentially private manner has also been studied; for example, see [14, 15, 16, 17, 18]. Generating a single sample in a private manner is the first step towards releasing a larger synthetic dataset, and to this end, we focus on the former challenge in this paper.
The remainder of this paper is organized as follows. We review various theoretical concepts relevant to the problem of differentially private sampling in Section II. In Section III, we present our proposed private sampling algorithm ROO and prove that it achieves $\epsilon$-DP with $\alpha$-accuracy. In Section IV, we present our generalized sampler DS-ROO, prove that it achieves $\epsilon$-DP, and provide upper bounds on its utility. Where relevant, we provide proof sketches of our theorems. We present numerical comparison of the utility of various private sampling algorithms in Section V and show that DS-ROO achieves superior performance for the same privacy guarantee, and conclude in Section VI.
@@ II Problem Setup
Definition 1 (Accuracy of Sampling).
An algorithm $\mathcal{A}$ is $\alpha$-accurate on a distribution $P$ if
@@
Definition 2 (Differential Privacy [2]).
A randomized algorithm, or mechanism $\mathcal{A}:\mathcal{X}^{n}\to\mathcal{Y}$ is considered $\epsilon$-differentially private ($\epsilon$-DP) if, for every pair of neighboring datasets $x^{n}\sim\tilde{x}^{n}\in\mathcal{X}^{n}$, and for $B\subseteq\mathcal{Y}$,
$\Pr\{\mathcal{A}(x^{n})\in B\}\leq e^{\epsilon}\Pr\{\mathcal{A}(\tilde{x}^{n})\in B\}.$ (4)
II-A State-of-the-art Algorithms
@@
  1. 1.
    Computes, for each $j\in[k]$, the empirical probability distribution $\hat{p}_{j}$,
  2. 2.
    Adds noise drawn from $\text{Lap}(\frac{2}{n\epsilon})$ to each count,
  3. 3.
    Uses an $L_{1}$ projection to restrict the noisy distribution to be a probability vector $\tilde{P}=(\tilde{p}_{1},\dots,\tilde{p}_{k})$, and
  4. 4.
    Outputs an element of $[k]$ sampled from $\tilde{P}$.
They show that this algorithm is $\alpha$-accurate with a sampling complexity
$n^{\prime}=\frac{2k}{\alpha\epsilon}.$ (5)
Cheu and Nayak [9] propose Subsampled Randomized Response (SubRR), which selects a data point uniformly at random from the input dataset and applies $k$-ary randomized response [20, 2]. In [9, Theorem 12], they show that their algorithm is $\alpha$-accurate with a reduced sampling complexity
$n^{\prime\prime}=\frac{1}{\alpha\epsilon}(k-1)(1-\alpha).$ (6)
In the following section, we prove that it is possible to achieve an $\epsilon$-DP and $\alpha$-accurate sampler with fewer samples than both $n^{\prime}$ and $n^{\prime\prime}$. Notably, for higher values of $\epsilon$, i.e., lower privacy, we gain an exponential reduction in the required number of samples. Moreover, [8] also establishes a lower bound on the sampling complexity as $\Omega\left(\frac{k}{\alpha\epsilon}\right)$ for a restricted range of $\epsilon\in(0,1]$.
@@ III Reveal-or-Obscure (ROO)
Algorithm 1 presents our proposed private sampler ROO. We implement the idea of obscuring the empirical distribution $\hat{P}_{x^{n}}$ by sampling from the uniform distribution on $[k]$. However, we wish to do so with a small probability $q$, so that we do not deviate too much from the true distribution $P$. With probability 1q1-q, we simply choose a sample from the given dataset, i.e., we reveal $\hat{P}_{x^{n}}$. We refer to $q$ as the obscuring probability, since it quantifies how often ROO obscures the empirical distribution by outputting a uniformly distributed sample instead.
Algorithm 1 Reveal-or-Obscure (ROO)
1:Dataset $x^{n}=(x_{1},\dots,x_{n})$, alphabet size $k$, privacy budget $\epsilon$, obscuring probability $q$
@@
3:With probability $q$, choose $y\sim\mathrm{Unif}[1\!:\!k]$
4:Otherwise, sample $i\sim\mathrm{Unif}[1\!:\!n]$ and set $y\leftarrow x_{i}$
5:return $y$
The privacy and utility guarantees provided by Algorithm 1 are given by the following theorem.
Theorem 1.
Given $q$, ROO (Algorithm 1) is $\epsilon$-DP and $\alpha$-accurate for
@@
Remark 1.
For a given dataset size $n$, alphabet size $k$ and privacy budget $\epsilon$, the corresponding obscuring probability $q$ can be obtained by rearranging the privacy expression (7) in Theorem 1:
$q=\frac{1}{1+\frac{n}{k}(e^{\epsilon}-1)}.$ (10)
@@
The proof consists of two parts—a privacy analysis, which establishes the $\epsilon$-DP guarantee, and a utility analysis, which derives the $\alpha$-accuracy bound.

### v3

# Reveal-or-Obscure: A Differentially Private Sampling Algorithm for Discrete Distributions

@@ Abstract
We introduce a differentially private (DP) algorithm called reveal-or-obscure (ROO) to generate a single representative sample from a dataset of $n$ observations drawn i.i.d. from an unknown discrete distribution $P$. Unlike methods that add explicit noise to the estimated empirical distribution, ROO achieves $\epsilon$-differential privacy by randomly choosing whether to “reveal" or “obscure" the empirical distribution. While ROO is structurally identical to Algorithm 1 proposed by Cheu and Nayak [1], we prove a strictly better bound on the sampling complexity than that extablished in Theorem 12 of [1]. To further improve the privacy-utility trade-off, we propose a novel generalized sampling algorithm called Data-Specific ROO (DS-ROO), where the probability of obscuring the empirical distribution of the dataset is chosen adaptively. We prove that DS-ROO satisfies $\epsilon$-DP, and provide empirical evidence that DS-ROO can achieve better utility under the same privacy budget of vanilla ROO.
@@ I Introduction
The task of synthetic data generation is closely related to the broader problem of learning probability distributions [4]. In the non-private setting, a learning algorithm approximates a distribution from which one can sample new data points that are representative of the original data. When privacy constraints are introduced, learning a distribution becomes significantly more challenging. In many practical cases, it may be sufficient to produce a small number of representative samples instead. The task of privately releasing one sample—known as DP sampling—is easier than full-fledged learning, since it requires less information from the underlying distribution. Motivated by this, we propose a novel DP sampling algorithm for discrete distributions on a finite alphabet. The key idea of our approach is to “obscure” the empirical distribution of the input dataset with a certain probability, or “reveal” it otherwise. Hence, we call our proposed algorithm reveal-or-obscure (ROO).
Main Contributions. Our main contributions are:
@@
  • We propose ROO—a sampling algorithm that achieves differential privacy without explicitly perturbing the empirical distribution of the input dataset. We incorporate uncertainty in our algorithm by sampling from the uniform distribution with some fixed probability $q$.
@@
  • We prove that our proposed algorithm reduces the sampling complexity while achieving better privacy-utility trade-off than the state-of-the art [5], [1].
@@
  • We also propose DS-ROO (data-specific ROO) as a technique to generalize ROO by making $q$, i.e., the probability of sampling from the uniform distribution, a function of the empirical distribution of the dataset. We prove that it is possible to achieve the same privacy guarantee with a lower $q$ value relative to the vanilla ROO algorithm for sufficiently large datasets.
@@
  • We demonstrate empirically that, for the same privacy guarantee, DS-ROO achieves better utility than vanilla ROO as well as the state-of-the-art in [5].
@@
Related Work. The problem of differentially private sampling from unknown distributions is first investigated in [5]. Raskhodnikova et al. [5] provide the first known bounds with $(\epsilon,\delta)$-DP guarantees on the complexity of sampling from arbitrary distributions over a discrete alphabet. DP algorithms for sampling from higher dimensional distributions such as multivariate Gaussians are presented in [6]. Husain et al. [7] considers DP sampling in the local setting, where the central aggregator cannot be trusted and each user must produce a single data record privately. Private sampling has also been studied in the distributed setting [8, 9]. A key focus of these efforts has been to reduce the sample complexity of DP-assured private sampling. While a recent and concurrent work [1] independent from ours proposes an algorithm structurally identical to ROO, our analysis establishes a strictly better sampling complexity bound in the same setting. More generally, the problem of releasing a dataset in a differentially private manner has also been studied; for example, see [10, 11, 12, 13, 14]. Generating a single sample in a private manner is the first step towards releasing a larger synthetic dataset, and to this end, we focus on the former challenge in this paper.
@@ II Problem Setup
Definition 1 (Accuracy of Sampling [15])
An algorithm $\mathcal{A}$ is $\alpha$-accurate on a distribution $P$ if the total variation distance, $d_{TV}$ between $Q$ and $P$ is bounded by some constant $\alpha$, i.e.,
@@
Definition 2 (Differential Privacy [2])
A randomized algorithm, or mechanism $\mathcal{A}:\mathcal{X}^{n}\to\mathcal{Y}$ is considered $\epsilon$-differentially private ($\epsilon$-DP) if, for every pair of neighboring datasets $x^{n}\sim\tilde{x}^{n}\in\mathcal{X}^{n}$, and for all $Y\subseteq\mathcal{Y}$,
$\Pr\{\mathcal{A}(x^{n})\in Y\}\leq e^{\epsilon}\Pr\{\mathcal{A}(\tilde{x}^{n})\in Y\}.$ (4)
@@
(i) computes, for each $j\in[k]$, the empirical probability distribution $\hat{p}_{j}$,
(ii) adds Laplace noise to each count,
(iii) uses an $L_{1}$ projection to restrict the Laplace-noised distribution to be a probability vector $\tilde{P}=(\tilde{p}_{1},\dots,\tilde{p}_{k})$, and
(iv) outputs an element of $[k]$ sampled from the distribution $\tilde{P}$.
For this algorithm, they show that for a sampling complexity
$n^{\prime}=\frac{2k}{\alpha\epsilon},$ (5)
their algorithm is $\alpha$-accurate. In the following section, we prove that it is possible to achieve an $\epsilon$-DP and $\alpha$-accurate sampler with fewer samples than $n^{\prime}$. Notably, for higher values of $\epsilon$, i.e., lower privacy, we gain an exponential reduction in the required number of samples. Moreover, [5] also establishes a lower bound on the sampling complexity as $\Omega\left(\frac{k}{\alpha\epsilon}\right)$ for a restricted range of $\epsilon\in(0,1]$.
@@ III Reveal-or-Obscure (ROO)
Algorithm 1 presents our proposed private sampler ROO. We implement the idea of obscuring the empirical distribution $\hat{P}_{x^{n}}$ by sampling from the uniform distribution on $[k]$. However, we wish to do so with a small probability $q$, so that we do not deviate too much from the true distribution $P$. With probability 1q1-q, we simply choose a sample from the given dataset, i.e., we reveal $\hat{P}_{x^{n}}$.
Algorithm 1: Reveal-or-Obscure (ROO)
1:Dataset $x^{n}=(x_{1},\dots,x_{n})$, alphabet size $k$, privacy budget $\epsilon$, parameter $q$
@@
3:With probability $q$, choose $y\sim\text{Unif}[1:k]$;
4:Otherwise, pick $i\sim\text{Unif}[1:n]$ and choose $y=x_{i}$;
5:return $y$;
The privacy and utility guarantees provided by Algorithm 1 is given by the following theorem.
Theorem 1
Given $q$, Algorithm 1 is $\epsilon$-DP and $\alpha$-accurate for
@@
Lemma 1
The proofs of Theorem 1 and Lemma 1 are in Appendix -A and -B, respectively.
In fact, the sampling complexity of Algorithm 1 is exponentially better in terms of $\epsilon$ compared to that of [5] in (5). This would appear to violate their lower bound on the sampling complexity; the reason it does not is that their lower bound only applies for $\epsilon<1$.
IV Data-Specific Reveal-or-Obscure (DS-ROO)
For the ROO algorithm in Section III, we fix $q$ in order to achieve $\epsilon$-DP for any possible dataset. From (27) in the privacy analysis of Appendix -A, we observe that the supremum of ratio of probabilities is achieved by setting $p=0$. This corresponds to the case where one of two neighboring datasets under consideration is entirely missing an element of the alphabet, but this element is present in the other dataset. Thus, ROO is inefficient on datasets where each element of the alphabet appears reasonably often. In this section, we show that by making $q$ a function of the dataset—specifically, a function of the smallest empirical probability, the accuracy can be improved for the same privacy. Let $m\in\left\{0,1,\dots,\left\lfloor\frac{n}{k}\right\rfloor\right\}$ denote the smallest number of times an element of $[k]$ appears in dataset $x^{n}$. Mathematically, $m$ can be expressed as $m=n\cdot\min_{x}\hat{P}_{x^{n}}(x)$. The modified private sampler—which we call the data-specific reveal-or-obscure (DS-ROO)—is presented in Algorithm 2. In DS-ROO, the probability $q_{m}$ is determined from the value of $m$ associated with the given dataset. Notably, when $m=0$, the corresponding $q_{0}$ is equivalent to that of Algorithm 1. That is, the case of a dataset that is missing an element of the alphabet (so $m=0$) is the worst-case scenario, and so in this case DS-ROO behaves identically to ROO. However, we will show empirically that in other cases, DS-ROO can do much better than vanilla ROO.
Algorithm 2: Data-specific ROO (DS-ROO)
2:Sample $y$
3:Compute
$q_{0}$ $=\frac{1}{1+\frac{n}{k}\left(e^{\epsilon}-1\right)};$ (9)
4:for $m=1,2,\dots,\left\lfloor\frac{n}{k}\right\rfloor$ do
5: Compute
$u_{m}$ $=-\frac{m}{n}+\frac{1}{k}-\frac{1}{n},$ (10)
$v_{m}$ $=e^{\epsilon}\left(\frac{1}{k}-\frac{m}{n}\right),$ (11)
$w_{m}$ $=-\frac{1}{n}-\frac{m}{n}+\frac{m}{n}e^{\epsilon},$ (12)
$q_{m}$ $=\max\left\{0,\frac{u_{m}}{v_{m}}q_{m-1}-\frac{w_{m}}{v_{m}}\right\};$ (13)
6:end for
7:Compute $m=n\cdot\displaystyle\min_{x}\hat{P}_{x^{n}}(x)$;
8:With probability $q_{m}$, choose $y\sim\text{Unif}[1:k]$;
9:Otherwise, pick $i\sim\text{Unif}[1:n]$ and choose $y=x_{i}$;
10:return $y$;
Fig. 1 shows the function $q_{m}$ for $k=9,n=1000$, and several different values of the privacy parameter $\epsilon$. We observe that $q_{m}$ is non-increasing in $m$. As $m$ increases, the empirical distribution of the dataset gets closer to the uniform distribution, and $q_{m}$ approaches zero. Thus, DS-ROO is less likely to obscure the empirical distribution for larger $m$. In high privacy regimes, $q_{m}$ tends to decrease much slower. On the other hand, for larger $\epsilon$, i.e., low privacy, $q_{m}$ goes to zero much faster. The privacy guarantee provided by DS-ROO is given by the following theorem.
Theorem 2
Algorithm 2 is $\epsilon$-differentially private.
The complete proof of Theorem 2 is provided in Appendix -C. We present a proof sketch below.
@@ IV Data-Specific Reveal-or-Obscure (DS-ROO)
Recall that $m$ denotes the smallest number of times an element of the alphabet $[k]$ appears in $x^{n}$. For neighboring datasets $x^{n}\sim\tilde{x}^{n}$, let $\tilde{m}$ denoted the $m$ value for $\tilde{x}^{n}$. Given $m$, there are three possible values of $\tilde{m}$: $(1)$ $\tilde{m}=m$, $(2)$ $\tilde{m}=m+1$, and $(3)$ $\tilde{m}=m-1$. Considering each of these cases separately and applying Definition 2, we derive the conditions under which DS-ROO satisfies $\epsilon$-DP as
$q_{m}\geq\frac{w_{m}}{u_{m}-v_{m}},\text{ for }m=0,1,\ldots,\left\lfloor\frac{1}{e^{\epsilon}-1}\right\rfloor,$ (14)
where $u_{m},v_{m}$, and $w_{m}$ are as defined in (10)–(12). Additionally, for $q_{m}$ to be a valid probability, we must also have $0\leq q_{m}\leq 1$ for all $m$. Setting the initial value $q_{0}$ as in (9), and assuming (16) to be an equality, we arrive at the expression
$q_{m}=\max\left\{0,\frac{u_{m}}{v_{m}}q_{m-1}-\frac{w_{m}}{v_{m}}\right\}.$ (17)
In Lemmas 3 and 4 of Appendix -D, we show that the above form of $q_{m}$ satisfies the inequality conditions in (14) and (15), respectively. Therefore, Algorithm 2 is $\epsilon$-differentially private. ∎
IV-A Utility of DS-ROO
We do not have theoretical bounds on the utility of DS-ROO at this time. However, we provide empirical evidence that DS-ROO achieves better utility than vanilla ROO and the state-of-the-art sampler in [5] for the same privacy guarantee. In order to measure the utility of DS-ROO, we consider an input distribution, estimate the corresponding output distribution according to Algorithm 2, and compute the total variation distance. Fig. 2 shows an example case for a distribution on an alphabet of size $k=9$, with dataset size $n=1000$, and privacy parameter $\epsilon=0.1$. We observe that mixing with the uniform distribution shifts some of the weight from the most probable central element to those with lower probabilities, resulting in reduced skewness in the output distribution.
Figure 2: Comparison of true distribution and the estimated output distribution for DS-ROO.
Figure 3: Comparison of accuracy ($\alpha$) versus privacy ($\epsilon$) curves of differentially private sampling algorithms.
Recall from Definition 1 that the total variation distance is upper bounded by $\alpha$. The smaller the value of $\alpha$, the more accurate the sampler is. From (5), the accuracy of the state-of-the-art sampler in [5] is
$\alpha=\frac{2k}{n^{\prime}\epsilon}.$ (18)
From Theorem 1, we have the accuracy of ROO
$\alpha=\frac{1}{1+\frac{n}{k}(e^{\epsilon}-1)}\left(1-\frac{1}{k}\right).$ (19)
For fixed values of $k$ and $n$, we obtain the accuracy of DS-ROO empirically, and compare it with that of [5] and ROO. Fig. 3 shows the $\alpha$ vs. $\epsilon$ curves for all three algorithms. We observe that DS-ROO achieves dramatically better accuracy than [5] and vanilla ROO while providing the same privacy guarantee. Of course, the numerical results in Fig. 3 are for the particular distribution shown in Fig. 2. We anticipate that we will see similar improvements for distributions that are not too skewed—if the distribution is more skewed (such as if the probability of a letter is 0), then there will be no improvement over vanilla ROO.
V Conclusion
In this work, we propose a novel differentially private sampling algorithm for discrete distributions on a finite alphabet. Our algorithm achieves differential privacy by obscuring the empirical distribution of a dataset without perturbing it directly. In addition, we propose a method to generalize our approach to achieve better utility for the same privacy guarantee. For future work, we aim to explore the practicality of our proposed algorithms for more complex distribution classes.
-A Proof of Theorem 1